According to the data listed at the Philippine Statistics Authority, the Philippine total


population reached 4 billion persons in year 2020 and was then increasing at the rate of 250


thousand persons each day. Assuming that natural population growth at this rate continues, what


will be the population at year 2060?

1
Expert's answer
2022-02-15T04:58:02-0500

4 billion = 2020 


Increase rate= 250,000 per day


2060=?


Population growth= x(t) = x0 × (1 + r) t


Initial Population (X0)=4000000000


Rate of Growth (%) =(r)9.125


# of years (t)=40


x(t)=40,000,000,000"\times" (1+9.125)40


x(t)=40,000,000,000"\times" 10.125"\times" 40


x(t)= 131531647490.14851


Population at year 2060= 131531647490 people.